# 0000-0002-5763-8174

### v1

@@ A Implementation Details
$\mathcal{L}_{\text{spatial}}(\hat{{\bm{\alpha}}},{\bm{\alpha}})=\sum_{v\in\{\text{WM},\text{GM},\text{CSF}\}}\lambda_{v}\,\lVert\hat{{\bm{\alpha}}}_{v}-{\bm{\alpha}}_{v}\rVert_{1},$ (4)
where ${\bm{\alpha}}_{v}$ and $\hat{{\bm{\alpha}}}_{v}$ denote the ground truth and estimated coefficient images restricted to tissue class $v$ (obtained from the segmentation masks computed from the GT reconstruction). We set $\lambda_{\text{WM}}=\lambda_{\text{GM}}=10$ and $\lambda_{\text{CSF}}=1$. This weighting reflects the fact that the acquisition protocol is specifically optimized for sensitivity to WM and GM, and we therefore prioritize accurate estimation in these tissue classes.
GLEAM stage

### v2

@@ A.2 Pretraining
$\mathcal{L}_{\text{spatial}}(\hat{{\bm{\alpha}}},{\bm{\alpha}})=\sum_{v\in\{\text{WM},\text{GM},\text{CSF},\text{Other}\}}\lambda_{v}\,\lVert\hat{{\bm{\alpha}}}_{v}-{\bm{\alpha}}_{v}\rVert_{1},$ (4)
where ${\bm{\alpha}}_{v}$ and $\hat{{\bm{\alpha}}}_{v}$ denote the ground truth and estimated coefficient images restricted to tissue class $v$ (obtained from the segmentation masks computed from the GT reconstruction), and ‘Other‘ corresponds to voxels outside the WM/GM/CSF masks. We set $\lambda_{\text{WM}}=\lambda_{\text{GM}}=10$ and $\lambda_{\text{CSF}}=\lambda_{\text{Other}}=1$. This weighting reflects the fact that the acquisition protocol is specifically optimized for sensitivity to WM and GM, and we therefore prioritize accurate estimation in these tissue classes.
A.3 GLEAM stage